OpenStudy (anonymous):

if f(x)=3x-2 and g(x)=6-4x, find f(x)+g(x) a. 7+4x b. 4-x c. -x-4 d. -4-7x

OpenStudy (anonymous):

f(x)+g(x)= 3x-2+6-4x =4-x

OpenStudy (anonymous):

thanks